OpenStudy (anonymous):

WHAT IS THE CONJECTURE BASED ON WHAT YOU OBSERVE. 11X11= 111X111= 1111X1111= 11111X11111=

OpenStudy (ybarrap):

for n = 1, (10^1 + 1)^2 for n = 2, (100 ^2+ 1)^2 for n = 3, (10^3 + 1)^2 for n = n, (10^n + 1)^2 Ans. For row n, factor will be (10^n + 1)^2